OpenStudy (anonymous):

Suppose that the graph of f contains the point (4,5). Find a point that must be on the graph of g(x)=2+f(6x) Similar prob but the multiplication is confusing me... (6 times x)

hero (hero):

Wait...lol

hero (hero):

it's 4/6, 7

OpenStudy (anonymous):

f(4)=5 so 6 times x = 4 but nothing times 6 would equal 4 that is my dilemma

hero (hero):

I can explain why, but check it first. That should be right

hero (hero):

6*4/6 = 4 Do you get it?

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

please explain

hero (hero):

I know it is. I knew we could figure this out

hero (hero):

Okay, so.... basically, we know that f(4) = 5 If you understand that much, you can figure this out. You also know that 6x must equal 4 so basically we have to find x. Which we found to be 6/4 so that was the hard part. now that we found x, we need to find g(x). Since we already know that f(4) = 5. All we have to do is replace f(6x) with 5: g(x) = 2 + f(6x), given (4,5) f(4) = 5 6x = 4 x = 4/6 g(x) = 2 + f(6x) g(x) = 2 + 5

hero (hero):

Mariochev, you good?

OpenStudy (anonymous):

Mariochev lol ...Yes I am sir. ..for now. haha. I'll let you know if I need more help.